For the year ended December 31, Depot Max’s cost of merchandise sold was $85,600.  Inventory at the beginning of the year was $8,000.  Ending inventory was $9,250.  Depot Max’s number of days' sales in inventory is closest to

Select one:

a.

34

b.

39

c.

74

d.

44
Answer #1

Average inventory = (8,000+9,250)/2 = 8625

Inventory Turnover = Cost of Goods Sold/Average inventory

= 85,600/8625 = 9.92 times

Days sales in inventory = 365/Inventory Turnover

= 365/9.92

= 37

Option B for S the nearest